# LIME: Less Is More for MLLM Evaluation

### v2

@@ A.2 Pipeline details
We chose GPT-4V as the basis for automatic judgment and made requests to the GPT-4V API using the prompt format below. For different subtask types, we used different prompt templates.
For VQA tasks, the prompt template is as follows:
Please judge whether the <Answer> is the golden answer to the <Question>.
If it is, please reply YES, otherwise reply NO.
<Question>: {question}
<Answer>: {answer}
<Your judgement>: <YES or NO>
For captioning tasks, the prompt template is as follows:
Now there is an image captioning task.
Please first describe the content of the image, then compare the image content with the provided captions.
If the captions are suitable as captions for the image, please answer YES; if they are not suitable, please answer NO.
Respond with NO if any of the captions are unsuitable. Respond with YES only if all captions are suitable.
You only need to reply with either YES or NO in <Your judgment>.
<Captions>: {answer}
<Description>: <Content of the image>
<Your judgement>: <ONLY YES or NO>
@@
[…]
A.2.2 metrics
Subtask metrics: As shown in the table 6, different metrics are used for different subtasks. It is important to note that, except for the CIDEr metric, all other metrics have a range between 0 and 1. The final score for each subtask is calculated by taking the average of these metrics.
Metric Subtask Formula Accuracy AI2D, ScienceQA-IMG, OCRBench, POPE CIDEr TextCaps,COCO-Caption Match score OK-VQA,TextVQA ANLS InfoVQA Relaxed Overall ChartQA Table 6: Metrics for different subtask -
@@
(1) (2) The arithmetic mean directly calculates the average of each subtask’s scores, while the weighted mean takes into account the number of samples in each subtask. We compare the results of these two calculation methods, as shown in the table 7. Except for Cogvlm, the scores calculated by the weighted average are lower than those by the arithmetic mean, and they show a higher correlation with WV-ELO. This suggests that the weighted average method is slightly superior to the arithmetic mean, as it considers the impact of the number of data points on the overall score, thereby avoiding potential errors caused by uneven data distribution. Therefore, in our work, we ultimately chose the weighted average as the method for calculating the overall score.
model overall_weighted overall_sum overall_cider WV_bench LLaVA-1.6-vicuna-7B 29.27 30.46 36.07 992 LLaVA-1.6-vicuna-13B 35.93 36.52 41.04 956 LLaVA-1.6-34B 42.73 43.30 47.12 1059 CogVLM 49.53 47.66 44.03 1016 Deepseek-VL 37.17 39.04 43.31 979 Idefics2 34.85 38.43 43.83 965 MiniCPM-v-1.0 25.58 28.95 35.79 910 Tinny-LLaVA-1-hf 13.76 17.79 24.15 879 LLaVA-1.5-13B 20.12 22.88 32.3 891 InstructBLIP-vicuna-7B 15.30 18.61 29.56 862 correlation score 0.91 0.90 0.87 1 Table 7: Comparison of different overall metrics method
A.2.3 difficulty classification details
For subtasks using the accuracy (acc) metric, where the scores are binary, with only 1 or 0, other tasks may have various possible score distributions (e.g., COCO-Caption, OK-VQA). Therefore, we determine the threshold score based on the overall distribution of subtask scores, and choose the cutoff value that offers the greatest distinction, as shown in table 8, for the metrics ANLS, Relaxed Overall and Accuracy (Acc), the threshold is set to 1.0, for BLEU-4 (for the captioning task, we use the BLEU-4 metric to represent the score for each question), the threshold is set to 0.2, while for Match Score, it is set to 0.6. When the score is greater than the threshold, it is marked as correct; otherwise, it is marked as incorrect.
@@
A.3 Alignment With Real World User Queries
To bridge the gap between LIME and real-world users’ query. We construct a similarity search system between the LIME and the real-world benchmarks(such as wildvision bench, and VibeEval) Different from single-modal similarity search systems, our system leverages multimodal information as query instructions, which contains text and image features information, to explore the effectiveness of the methods and construct the optimal similarity search system, we propose the following different approaches:
single modal query
For single-modal query, we use the corresponding modality information for retrieval, such as text-to-text and image-to-image. The CLIP model, which consists of a text encoder and an image encoder (Radford et al., 2021), has achieved excellent performance in representing multimodal information. As shown in Figure 9, We use CLIP’s encoders to encode single-modal features and use cosine similarity as measure metrics. For each data entry in the real-world benchmark, we index the top-k most similar entries sequentially. -
multi-modal query For multimodal queries, we combine image features and text features as the query. As shown in Figure 9, We adapt both the Clip and Blip models to combine these features. For the clip model, the text and image inputs are processed into separate vector features, which are then combined into a unified vector through a weighted sum. As to the blip model, image features are fused with text features through a cross-attention mechanism. We only keep the top-1 similar data entry based on similarity.
In Figure 10, we calculate the correlation coefficient between LIME-fit obtained through different query methods and real-world benchmarks. We find that samples retrieved using image queries align better with the WildVision benchmark compared to those retrieved with text queries. In multimodal VQA tasks, text queries often suffer from various forms of repetition, whereas images are unique and provide more significant information.
However, whether using a single-modal or multi-model query method, the Correlation Coefficient obtained by all methods is lower than the baseline score, which suggests that directly querying results from the origin data does not align well with real-world user queries.
Why current multimodal benchmarks can’t reflect real user queries in the wild real world?
Instructions in original benchmark can not address real-world problems. Although there are various multimodal evaluation datasets (TextVQA, infoVQA, ScienceQA, and MMMU) , the instructions in these datasets differs from real-world user instructions. For the same image, the types of instructions can also vary significantly. As illustrated in Figure11, typical VQA tasks require perception and reasoning about detailed content within an image, such as asking ’the population older than 65 in Vietnam in 2020’. In contrast, real-world queries might ask ’how to reproduce the charts using Python’. This discrepancy leads to inherent differences in instructions, even when using retrieval algorithms to find the most relevant dataset questions.
Back to List of figures
Back to List of figures
Back to List of figures
Back to List of figures
Back to List of figures
Back to List of figures
Back to List of figures
Back to List of figures
Back to List of figures
Back to List of figures
Back to List of figures
Back to List of figures
Back to List of figures
Back to List of figures
Back to List of figures
Back to List of figures

### v3

@@ A.3 Pipeline details
We selected GPT-4V as the basis for automatic judgment and interacted with the GPT-4V API using specific prompt templates for different subtasks.
@@
A.3.2 metrics
Subtask metrics: As shown in the Tab 7, different metrics are used for different subtasks. It is important to note that, except for the CIDEr metric, all other metrics have a range between 0 and 1. The final score for each subtask is calculated by taking the average of these metrics.
| Metric | Subtask | Formula |
| Accuracy | AI2D, ScienceQA-IMG, OCRBench, POPE | |
| CIDEr | TextCaps,COCO-Caption | |
| Match score | OK-VQA,TextVQA | |
| ANLS | InfoVQA | |
| Relaxed Overall | ChartQA |
@@
The arithmetic mean directly calculates the average of each subtask’s scores, while the weighted mean takes into account the number of samples in each subtask. We compare the results of these two calculation methods, as shown in the Tab 8. weighted average method achieves a higher correlation with WV-ELO. This suggests that the weighted average method is slightly superior to the arithmetic mean, as it considers the impact of the number of data points on the overall score, thereby avoiding potential errors caused by uneven data distribution. Therefore, in our work, we ultimately chose the weighted average as the method for calculating the overall score.
| model | overall_weighted | overall_sum | overall_cider | WV_bench |
| LLaVA-1.6-vicuna-7B | 30.15 | 30.46 | 36.07 | 992 |
| LLaVA-1.6-vicuna-13B | 37.08 | 36.52 | 41.04 | 956 |
| LLaVA-1.6-34B | 44.06 | 43.30 | 47.12 | 1059 |
| CogVLM | 51.03 | 47.66 | 44.03 | 1016 |
| Deepseek-VL | 38.1 | 39.04 | 43.31 | 979 |
| Idefics2 | 36.39 | 38.43 | 43.83 | 965 |
| MiniCPM-v-1.0 | 26.15 | 28.95 | 35.79 | 910 |
| Tinny-LLaVA-1-hf | 13.95 | 17.79 | 24.15 | 879 |
| LLaVA-1.5-13B | 20.38 | 22.88 | 32.3 | 891 |
| InstructBLIP-vicuna-7B | 15.55 | 18.61 | 29.56 | 862 |
| correlation score | 0.91 | 0.90 | 0.87 | 1 |
A.3.3 difficulty classification details
For subtasks using the accuracy (acc) metric, where the scores are binary, with only 1 or 0, other tasks may have various possible score distributions (e.g., COCO-Caption, OK-VQA). Therefore, we determine the threshold score based on the overall distribution of subtask scores, and choose the cutoff value that offers the greatest distinction, as shown in Tab 9, for the metrics ANLS, Relaxed Overall and Accuracy (Acc), the threshold is set to 1.0, for BLEU-4 (for the captioning task, we use the BLEU-4 metric to represent the score for each question), the threshold is set to 0.2, while for Match Score, it is set to 0.6. When the score is greater than the threshold, it is marked as correct; otherwise, it is marked as incorrect.
@@
A.3.4 Retrieve from real world queryd
Appendix B case study
The original dataset contains noise data. In the following figure, we categorize the problematic data into three types and present specific examples from different datasets.
Text Answerable Questions:
Some questions can be answered without the need for visual information, mainly focusing on the AI2D and ScienceQA datasets. As shown in figs. 11 and 12, AI2D and ScienceQA emphasize knowledge in the field of science while overlooking the importance of visual information. Given the background of domain knowledge, some LLMs are able to provide answers even without requiring visual input.
Annotation Error Questions:
Repeated Question:
Some benchmarks also contain a significant amount of duplicate data, where the question content and image content are completely identical. This issue is mainly found in the POPE dataset, as shown in the figs. 15, 16, 17, 18 and 19.